OpenStudy (anonymous):

Find the exact value of the area between the graph of y = cosx and y = e^x for 0 ≤ x ≤ 1.

OpenStudy (anonymous):

i know that the antiderivative of cos(x) is sin(x) and the antiderivative of e^x is e^x and since the interval is from 0 to 1 i know that i would do cos(1)-cos(0)= some value and then e^1-e^0=some value, but its asking for an EXACT answer so idk what to put. I write down a decimal and it says its wrong :/

myininaya (myininaya):

ok so we do \[\int\limits_{0}^{1}(e^x-\cos(x))dx \]

myininaya (myininaya):

why do you plug in 1 in limits into the cos function

myininaya (myininaya):

you plug in limits into the antiderivative

myininaya (myininaya):

not the original function

OpenStudy (anonymous):

oh yeah sorry thats my mistake i meant sin (x)

OpenStudy (anonymous):

so its e^x-sinx

myininaya (myininaya):

so what do you get after integrating that e^x-sin(x) is correct so plug in your limits

OpenStudy (anonymous):

so i plug in e^1-sin(0)=2.718281828

myininaya (myininaya):

(e^1-sin(1))-(e^0-sin(0)) e^1-sin(1)-1+0 e-sin(1)-1<---this is exact answer

OpenStudy (anonymous):

ohhhh i see what i was doing wrong

OpenStudy (anonymous):

thank you!!!

myininaya (myininaya):

you were letting e^0 =0?

myininaya (myininaya):

a^0 =1 excluding when a=0

OpenStudy (anonymous):

yeah idk what i was doing; i always make those kinds of mistakes